OpenStudy (anonymous):

Will award medal and fan simplify ((x^2+y^2)^(3/2))^2

OpenStudy (anonymous):

x6 × x5 = x11

OpenStudy (danjs):

\[[(x^2 + y^2)^{3/2}]^2 \]

OpenStudy (anonymous):

yee

OpenStudy (anonymous):

x^6+3x^4 y^2+ 3 x^2 y^4 + y^6

OpenStudy (anonymous):

but not in expanded form

OpenStudy (danjs):

just (x^2 + y^2)(x^2 +y^2)

OpenStudy (danjs):

3 times

OpenStudy (anonymous):

I am not sure, that is what I know

OpenStudy (anonymous):

yea but i got this answer for my calc 2 review

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

because for the quotient rule you need to square the denominator, but i dont see where the 5/2 comes from

OpenStudy (danjs):

\[[x ^{3/2}]^2 = [x^3]\]

OpenStudy (anonymous):

i am not sure

OpenStudy (anonymous):

yea, i know thats what im saying.

OpenStudy (anonymous):

\[\left[ x^3/2 \right]^2 = \left[ x^3 \right]\]

OpenStudy (anonymous):

medal and fan

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

yea but whats up with the answer that i found online

OpenStudy (anonymous):

it is wrong!

OpenStudy (anonymous):

it did everything right for the partial integration

OpenStudy (anonymous):

my answers are always right medal plz

OpenStudy (anonymous):

although your mad pompous i did say id award u a medal